-find the reaction forces at A & B.
​-shear and moment diagrams
​-locate where the shear is zero and bending moments are at a maximum positive and negative. Convert any decimals to closest feet and inches 1/8 inch.

2.0m (o. o KN 10,0 KN 2.0m 2.om
Answer #1

Assumed A left side B right side

First calculate reaction at A and B

TOTAL UDL load, the load will be 4*10=40KN AND CENTRE LINE OF THE UDL IS 5 M FROM LEFT END

Now taking moment about B,

RA*8=40*5+10*2-6*

RA=26 KN

Now,

Now consider all upward forces are equal to downward forces for calculating RB

RA+RB=10+6+40=56

RB=56-26=30 KN

NOW sfd AND BMD can be drawn